A 5% solution of cane sugar is isotonic with 10% of a solution of an unknown solute the molar mass of unknown solute in g/mol

Grade:12th pass

1 Answers

Arun
6 years ago
Molecular mass of sugar cane= 342( it should be known)
Hence
Molecular mass of sugarcane/ % solution = molecular mass of unknown/ % solution
Hence
342/5 = x /10
X = 684
Hope it helps